# LV-UNet: A Lightweight and Vanilla Model for Medical Image Segmentation

### v2

@@ I Introduction
UNet  [4] is a classic medical image segmentation model consisting of an encoder and a decoder with skip connections, showing a significant improvement compared to previous attempts. UNet’s variants like Attention-UNet  [5], U-Net++  [6], ResUnet  [7], Unet 3+ [8] and so on have achieved great success. Due to the impressive performance of Transformers in visual tasks in recent years, Transformer-based U-Net variants like Trans-UNet [9], Swin-UNet [10], MT-UNet [11] and Unetr [12] and so on have gained wider recognition and have shown remarkable effectiveness. However, these models entail significant computational overhead and a large number of parameters, making them difficult to use in point-of-care applications. Thus, more and more effort has been paid to designing lightweight models in medical image segmentation scenarios. Although existing lightweight medical image segmentation models have made significant progress, they often lack robustness across different datasets. How to address these issues by designing a model with reduced computational overhead, fewer parameters, faster inference time, and maintained performance, while also ensuring robustness across various datasets remains a challenge.
@@ IV-A Overall Architecture
The model architecture of LV-UNet contains an encoder, a decoder, and final output layers, which can be seen in Figure 1. The encoder includes pre-trained blocks from MobileNetv3-Large, as well as fusible encoder modules. The decoder consists of fusible decoder modules. There are skip connections between the corresponding layers in the encoder and decoder, and to avoid significantly increasing the number of parameters, element-wise addition is used instead of concatenation. The dimensions of each stage are 16, 24, 40, 80, 160, 240 and 480.
@@ IV-C1 Architecture
The training mode of a fusible encoder block contains a convolutional layer using a $1\times 1$ kernel with stride 1, a batch normalization operation, a Leaky ReLU function, a convolutional layer using a $1\times 1$ kernel, a pooling layer and a designed activation layer. The resolutions of output features of an encoder block are half of the input. In the deployment mode, the block only contains a convolutional layer, a pooling layer, and the designed activation layer. A fusible decoder block is similar, where the pooling operation is replaced by an upsampling operation. The resolutions of output features of a decoder block are two times of the input. The architectures of the fusible blocks in the training and deployment modes are shown in Figure 2. The method to transfer the training mode to the deployment mode during the inference stage will be introduced in the section ”Re-parametrization and Deployment Mode”.
@@ IV-C2 Nonlinear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. The nonlinear activation function is followed by a convolution and a batch normalization operation. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and nonlinear relationships within the input data.
@@ IV-C3 Re-parametrization and Deployment Mode
VanillaNet [21] adopts a deep training strategy in the training process, and it is also applied in the LV-UNet training. At the start of the training process, two convolutional layers utilize an activation function. During the training epochs, the activation function is progressively adjusted to resemble an identity mapping. At the end of the training, the two convolutions can seamlessly combine into a single convolution, resulting in reduced inference time. We use a Leaky ReLU function to control it:
@@ V-A1 ISIC2016
ISIC 2016 [13] is a dataset of dermoscopic images of skin lesions for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2016. In this paper, we used the lesion segmentation dataset, which contains original images and binary lesion masks. The training set has 900 pairs of original images and corresponding masks. In the testing set, there are 379 pairs of images and masks. The training set is split into a new training set and a validation set with a ratio of 8:2.
@@ V-A2 BUSI
Breast Ultrasound Images Dataset (BUSI) [14] is a dataset of breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, and corresponding masks. Unlike the setting in [18], we include all three classes in the dataset, and the test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [26]. The initial learning rate is 0.001, and a CosineAnnealingLR [27] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metrics in validation are $IOU$ of the lesions. The loss function used is a mixed loss that combined binary cross-entropy (BCE) loss and dice loss [28]:
@@
We also train, validate and test UNet, Attention UNet, UNet++, UNeXt, MLUNet, EGE-UNet, U-KAN and MobileNetv3-UNet (UNet with MobileNetv3 as the backbone) for comparison. UNet uses double convolution layers in each stage. The dimensions of each stage in UNet, Attention UNet and UNet++ are 32, 64, 128, 256 and 512. UNeXt, MLUNet, EGE-UNet, and U-KAN use the original design in their papers. MobileNetv3-UNet adds the output of each encoder stage to the blocks in the decoder.
For other experiments, we select the model that performs the best on the validation set for inference on the test set. However, the Leaky ReLU only transitions to an identity mapping function at the end of the training. Therefore, we choose the model from the last epoch for inference on the test set, for the LV-UNet using the deep training strategy.
@@ V-C Results
The comparison experimental results can be seen in Table \@slowromancapi@, from which we can see LV-UNet shows competitive performances compared to the baselines. The deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and FLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point-of-care scenarios. Visualizations of some segmentation results can be seen in Figure 3.
@@
UNet 7.8M 12.15G
Attention-UNet 8.7M 16.79G
UNet++ 9.2M 34.77G
UNeXt 1.5M 0.58G
MALUNet 0.2M 0.09G
EGEUNet 53.4K 0.08G
U-KAN 25.4M 10.5G
MobileNetv3-Large-UNet 2.0M 0.51G
@@ V-D Ablation Studies
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and FLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting (combination \@slowromancapii@) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination \@slowromancapi@, combination \@slowromancapii@ reduces the model size and computational complexities significantly.
The Number of Series-Informed Activation Functions. In experiments in [21], the increase of the number of series-informed activation functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. However, the results suggest that using an appropriate number of Series-Informed Activation Functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
TABLE VI: Comparison of adding and concatenation methods for Skip-Connection
@@
The Way of Skip-Connections. We evaluated the two kinds of Skip-Connections methods: adding and concatenation. The adding operation in Skip-Connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies. We also compared different nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in  [21] $a={e}/{E}$. The comparison results in Table \@slowromancapvii@ show that performances are very close and our method achieves better results on more datasets.

### v3

@@ I Introduction
UNet  [4] is a classic medical image segmentation model consisting of an encoder and a decoder with skip-connections, showing a significant improvement compared to previous attempts. UNet’s variants like Attention-UNet  [5], U-Net++  [6], ResUnet  [7], Unet 3+ [8] and so on have achieved great success. Due to the impressive performance of Transformers in visual tasks in recent years, Transformer-based U-Net variants like Trans-UNet [9], Swin-UNet [10], MT-UNet [11] and Unetr [12] and so on have gained wider recognition and have shown remarkable effectiveness. However, these models entail significant computational overhead and a large number of parameters, making them difficult to use in point-of-care applications. Thus, more and more effort has been paid to designing lightweight models in medical image segmentation scenarios. Although existing lightweight medical image segmentation models have made significant progress, they often lack robustness across different datasets. How to address these issues by designing a model with reduced computational overhead, fewer parameters, faster inference time, and maintained performance, while also ensuring robustness across various datasets remains a challenge.
@@ IV-A Overall Architecture
The model architecture of LV-UNet contains an encoder, a decoder, and final output layers, which can be seen in Fig.1. The encoder includes pre-trained blocks from MobileNetv3-Large, as well as fusible encoder modules. The decoder consists of fusible decoder modules. There are skip-connections between the corresponding layers in the encoder and decoder, and to avoid significantly increasing the number of parameters, element-wise addition is used instead of concatenation. The number of channels in each stage is 16, 24, 40, 80, 160, 240 and 480.
@@ IV-C1 Architecture
The training mode of a fusible encoder block contains a convolutional layer using a $1\times 1$ kernel with stride 1, a batch normalization operation, a Leaky ReLU function, a convolutional layer using a $1\times 1$ kernel, a pooling layer and a designed activation layer. The resolutions of output features of an encoder block are half of the input. In the deployment mode, the block only contains a convolutional layer, a pooling layer, and the designed activation layer. A fusible decoder block is similar, where an upsampling operation replaces the pooling operation. The resolutions of output features of a decoder block are two times of the input. The architectures of the fusible blocks in the training and deployment modes are shown in Fig.2. The method to transfer the training mode to the deployment mode during the inference stage will be introduced in the section ”Re-parametrization and Deployment Mode”.
@@ IV-C2 Nonlinear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and nonlinear relationships within the input data.
@@ IV-C3 Re-parametrization and Deployment Mode
VanillaNet [21] adopts a deep training strategy in the training process, and it is also applied in the LV-UNet training. At the start of the training process, two convolutional layers utilize an activation function. During the training epochs, the activation function is progressively adjusted to resemble an identity mapping. At the end of the training, the two convolutions can seamlessly combine into a single convolution, resulting in reduced inference time. We use a Leaky ReLU function between two convolutional layers to control it:
@@ V-A1 ISIC2016
ISIC 2016 [13] is a dataset of skin lesions images for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2016. In this paper, we used the lesion segmentation dataset, which contains original images and binary lesion masks. The training set has 900 pairs of original images and corresponding masks. In the testing set, there are 379 pairs of images and masks. The training set is split into a new training set and a validation set with a ratio of 8:2.
@@ V-A2 BUSI
Breast Ultrasound Images Dataset (BUSI) [14] is a dataset of breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, with corresponding masks. Unlike the setting in [18], we include all three classes in the dataset, and the test set is split with a ratio of 0.2. Then the rest of the dataset is split into the training set and the validation set with a ratio of 8:2.
@@ V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [26]. The initial learning rate is 0.001, and a CosineAnnealingLR [27] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metric in the validation is $IOU$. The loss function used is a mixed loss that combined binary cross-entropy (BCE) loss and dice loss [28]:
@@
We also train, validate and test several models like UNet  [4], Attention UNet  [5], UNet++ [6], UNeXt [18], MALUNet [19], EGE-UNet [20], U-KAN [25] and MobileNetv3-UNet (UNet with MobileNetv3 as the backbone) for comparison. UNet uses double convolution layers in each stage. The dimensions of each stage in UNet, Attention UNet and UNet++ are 32, 64, 128, 256 and 512. UNeXt, MLUNet, EGE-UNet, and U-KAN use the original design in their papers. MobileNetv3-UNet adds the output of each encoder stage to the blocks in the decoder.
For other experiments, we select the model that performs the best in the validation set for inference on the test set. However, the Leaky ReLU only transitions to an identity mapping function at the end of the training. Therefore, we choose the model from the last epoch for inference on the test set, for the LV-UNet using the deep training strategy.
@@ V-C Results
The experimental results of the comparison can be seen in Table \@slowromancapi@, using $IOU$ and $Dice$ as metrics, from which we can see that LV-UNet shows competitive performances compared to baselines. In addition, the deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and FLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point of care scenarios. Visualizations of some segmentation results can be seen in Fig.3.
@@
UNet [4] 7.8M 12.15G
Attention-UNet [5] 8.7M 16.79G
UNet++ [6] 9.2M 34.77G
UNeXt [18] 1.5M 0.58G
MALUNet [19] 0.2M 0.09G
EGEUNet [20] 53.4K 0.08G
U-KAN [25] 25.4M 10.5G
MobileNetv3-Large-UNet [22] 2.0M 0.51G
@@ V-D Ablation Studies
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and FLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting (combination \@slowromancapii@) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination \@slowromancapi@, combination \@slowromancapii@ reduces the model size and computational complexities.
The Number of Series-informed Activation Functions. In experiments in [21], the increase of the number of series-informed activation functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. However, the results suggest that the use of an appropriate number of series-informed activation functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
TABLE VI: Comparison of Adding and Concatenation Methods for Skip-Connections
@@
The Way of Skip-Connections. We evaluated the two types of skip-connections methods: adding and concatenation. The adding operation in skip-connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies. We also compared different nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in  [21] $a={e}/{E}$. The comparison results in Table \@slowromancapvii@ show that the performances are very close and that our method achieves better results in more datasets.
@@ VI Conclusions
Limitations The learning rate in our setting might not be optimal for larger models, and experiments were not repeated multiple times. Besides, utilizing a more advanced pre-trained model could be a promising direction, as MobileNetv3-Large is no longer the state-of-the-art lightweight model.